OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical (-1, -5) and (-4, -5)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Plug it int he slope formula

OpenStudy (igreen):

\(m = \dfrac{-5+5}{-4+1}\)

OpenStudy (anonymous):

-5 + 5 = 0, and -4 + 1 = -3 right?

OpenStudy (igreen):

Yep, so we have: 0 / -3 And anything divided from 0 is 0, so the slope is 0. And the line will be horizontal.

OpenStudy (anonymous):

Sweet! Your amazing. Thank you so much for helping me.